OpenStudy (anonymous):

What is the simplified radical of 64

OpenStudy (zale101):

8 * 8=64 8^2=64

OpenStudy (anonymous):

8

OpenStudy (anonymous):

8*8 = 64

OpenStudy (zale101):

\[\sqrt{64}=8\]

OpenStudy (anonymous):

thank u so much